OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 4) and whose slope is 1/2.

OpenStudy (accessdenied):

Do you know the "Point-Slope form"?

OpenStudy (anonymous):

No.

OpenStudy (accessdenied):

Hmm... Slope-intercept form, then? \(y = mx + b\)

OpenStudy (anonymous):

y=mx+b ,m=1/2 (2,4) 4=2/2+b ->b=3 thus y=x/2+3